A ball (ball A) was thrown with an initial velocity of 15 meters per second. After 1.15 seconds, another ball (Ball B) is thrown with an initial velocity of 12 m/s. At what height do they collide?


1
Expert's answer
2022-01-18T15:13:00-0500

"h(t)=vt-\frac{gt^2}2,"

"v_1 t-\frac{gt^2}2=v_2(t-t_0)-\frac{g(t-t_0)^2}2,"

"t=2.45~s,"

"h(2.45)=7.3~m."